OpenStudy (anonymous):

the point (6,2) lies on the line -x-8y=-22 determinme if this point is a solution of the system -x-8y=-22, -7x+3y=-36

OpenStudy (anonymous):

As long as it satisfies both equations, you can consider it to be a solution of the system, so since you know it is a solution to the first one, you just need to check the second. -7(6)+3(2)=-36 -42+6=-36 So since the point falls on both equations, you can say that it IS a solution of the system.